OpenStudy (anonymous):

the diameter of a baseball is 2.92 cm and the diameter of a softball is 3.84 cm. how many times the surface area of a softball is to a surface area of a baseball? round your answer to the nearest thousandth.

OpenStudy (anonymous):

the surface areas of the sphere is proportional to the square of the diameter

OpenStudy (anonymous):

hence reqd ans is (3.84/2.92)^2 = (1.31506)^2 =1.729 appx